# The geometry of Deroin–Tholozan representations

### v2

@@ 4.4.2. Injectivity of the realization map
The value of $\beta_{b}$ is determined, up to sign, by (4), and its sign is determined by the sign of $\kappa_{b}$:
@@
Next, we prove that $[\mathcal{F}]$ also determines the values of the slit angles. We’re assuming that the underlying pants assembly $\Upsilon$ is chained, so there are never more than two slits on a given samosa. That means we can find expressions for all the slit angles using Lemma 3:
@@ 4.4.8. Continuity of the inverse map
We now prove that the inverse map of the realization map, defined on $R(\mathcal{S}_{\alpha,\Upsilon}^{\varepsilon})\subset\operatorname{Hyp}_{\alpha}(S_{\Upsilon})$, is also continuous. To do so, we show that the samosa assembly parameters $(\beta_{b},\ell_{b},\phi_{b},\phi^{\prime}_{b})_{b\in{\mathcal{B}}}$ of $R^{-1}([\mathcal{F}])$ depend continuously on $[\mathcal{F}]\in R(\mathcal{S}_{\alpha,\Upsilon}^{\varepsilon})$. We can relate $\ell_{b}$ to $c_{b}$ and $\kappa_{b}$ by (10). Proposition 1 says that $c_{b}$ varies continuously in $[\mathcal{F}]$ and so does the kink angle $\kappa_{b}$. This shows that $\ell_{b}$ depends continuously on $[\mathcal{F}]$ for every $b\in{\mathcal{B}}$. The value of $\beta_{b}$ is expressed by (11) continuously in terms of $c_{b}$ and $\kappa_{b}$ showing that $\beta_{b}$ varies continuously in $[\mathcal{F}]$.
Finally, to prove that the slit angles $\phi_{b}$ and $\phi_{b}^{\prime}$ depend continuously on $[\mathcal{F}]$, we use (12) and (13). These equations express the slit angles as continuous functions of $\alpha,\beta,\ell,\lambda$, and, sometimes, $\xi$ and $\eta$. We already established the continuous dependence of $\ell_{b}$ and $\beta_{b}$ on $[\mathcal{F}]$. From Proposition 1 we deduce that both $\lambda$ and $\xi$ depend continuously on $[\mathcal{F}]$, and so does $\eta$. This finishes the proof that $R^{-1}$ is continuous.
@@ 4.4.9. Smoothness of the realization map and its inverse
It remains to prove that the realization map is a diffeomorphism onto its image under the assumption that Conjecture 1 holds. The argument is similar to the proof that the realization map is continuous. Using Conjecture 1, it’s enough to check that a finite number of arc lengths depend smoothly on the samosa parameters to prove that $R$ is smooth. We use (14), (15), (16), and (17) for that. Recall that we’re working on the domain $\mathcal{S}_{\alpha,\Upsilon}^{\varepsilon}$ which prescribes a hemisphere for each slit, and hence a sign for each slit angle. Similarly, to prove that the inverse map is smooth, it is sufficient to express the samosa parameters as smooth functions of the same finite set of arc lengths. Here, we use (10), (11), (12), and (13). For instance, from (11) we see that $\beta_{b}$ depends smoothly on the parameters $\kappa_{b}$ and $c_{b}$ except maybe when $\kappa_{b}=0$ or when
@@
It is however not hard to see that $\sin(\kappa_{b}/2)^{2}\cosh(c_{b})-\cos(\kappa_{b}/2)^{2}\geq-1$ with equality if and only if $\kappa_{b}=0$. Alternatively, it is also possible to write
@@
using the hyperbolic law of sines—formula (S) from Appendix A. This alternative expression of $\beta_{b}$ shows that it depends smoothly on $\kappa_{b}$, $c_{b}$ and $\ell_{b}$ around $\kappa_{b}=0$ as well, because $c_{b}$ and $\ell_{b}$ stay bounded away from 0 in this region. The treatment of the other variables is similar and we won’t provide more details here.
@@ 5.2. Unfolding samosas
A samosa with two slits in the same hemisphere is a hamantash as long as its slit angles satisfy $\phi_{a}\neq 0$ and $\phi_{b}^{\prime}\neq\pm\beta_{b}^{\prime}/2$, and its slit lengths $\ell_{a}$ and $\ell_{b}$ don’t exceed the respective bounds $\ell_{a,max}$ and $\ell_{b,max}$ given by the formulae
$\coth(\ell_{a,max})$ $=\frac{1}{\sinh(d)}\big{(}\cos(\phi_{a})\cosh(d)+|\sin(\phi_{a})|\cot(\beta_{b}^{\prime}/2-|\phi_{b}^{\prime}|)\big{)}$
$\coth(\ell_{b,max})$ $=\frac{1}{\sinh(d)}\big{(}\cos(\beta_{b}^{\prime}/2-|\phi_{b}^{\prime}|)\cosh(d)+\sin(\beta_{b}^{\prime}/2-|\phi_{b}^{\prime}|)\cot(|\phi_{a}|)\big{)},$
@@ 5.3. Unfolding samosa assemblies
The subspace of hamantash assemblies is denoted $\mathcal{H}_{\alpha,\Upsilon}\subset\mathcal{S}_{\alpha,\Upsilon}$. A hamantash assembly can be unfolded to a collection of hyperbolic polygons embedded in the hyperbolic plane. Using side identifications, all the pieces can be glued back together to produce the realization of the hamantash assembly. In other words, the process of gluing the hamantashen along their slits commutes with the processes of unfolding and re-folding the hamantashen.
When none of the slits lie on equator segments opposite from un-slit corners, as discussed in Section 4.3.4, the realized hamantash assembly comes with a canonical isotopy class of homeorphisms to the generic realization $S_{\Upsilon}$. Thus, the process of unfolding and then gluing is the restriction $R\colon\mathcal{H}_{\alpha,\Upsilon}^{\text{gen}}\to\operatorname{Hyp}_{\alpha}(S_{\Upsilon})$ of the realization map, where $\mathcal{H}_{\alpha,\Upsilon}^{\text{gen}}=\mathcal{S}_{\alpha,\Upsilon}^{\text{gen}}\cap\mathcal{H}_{\alpha,\Upsilon}$.
@@ 6.1. The holonomies of hamantashen
Given a chained pants assembly $\Upsilon$, let $\mathcal{H}_{\alpha,\Upsilon}^{\textrm{north}}\subset\mathcal{H}_{\alpha,\Upsilon}$ be the subset consisting of hamantash assemblies whose slits all lie in northern hemispheres. The realization map extends continuously from $\mathcal{H}_{\alpha,\Upsilon}^{\text{gen}}$ to $\mathcal{H}_{\alpha,\Upsilon}^{\text{north}}$. To see why, first recall from Section 4.3.4 why the realization map fails to extend from $\mathcal{S}_{\alpha,\Upsilon}^{\text{gen}}$ to $\mathcal{S}_{\alpha,\Upsilon}$. The problem is that a non-generic samosa assembly comes with several isometry classes of identifications with $S_{\Upsilon}$, and the choice between them can’t be made continuously throughout $\mathcal{S}_{\alpha,\Upsilon}$. In $\mathcal{H}_{\alpha,\Upsilon}^{\text{north}}$, we can make a continuous choice of identification by pushing all the slits infinitesimally into the northern hemisphere.
@@
can be understood as follows. Unfolding a hamantash assembly produces a collection of hyperbolic polygons embedded in the hyperbolic plane. We attach the polygons at each vertex coming from a pants curve $b\in\mathcal{B}$ by pairwise gluing all the sides corresponding to slits. An unfolded hamantash in $\mathcal{H}_{\alpha,\Upsilon}^{\textrm{north}}$ consists of a triangle, which is the southern hemisphere, and one to three flaps, which are the pieces of the northern hemisphere. Erasing the flaps gives a triangle chain, laid out along the spine of the hamantash assembly, with the geometric features described in Section 2.2. Since all the triangles in the chain are southern hemispheres, they all lie on the same side of the spine introduced in Section 4.3.4.
@@ 6.2. Finding a suitable pants decomposition
    For any pair of pants decomposition and any DT representation, there is always at least one non-degenerate triangle in the associated chain. This follows from the fact that the sum of all areas of the triangles in the chain is a positive constant (essentially equal to the Toledo number of the representation).
@@
Assume that $n\geq 4$, since the claim is obvious for $n=3$. For concreteness, identify $\Sigma$ with the one-point compactification of the plane, with the marked points laid out around a circle. We label the marked points $1,\ldots,n$ in counterclockwise fashion. The marked points then form a cyclically ordered set. Introduce a standard geometric presentation of $\pi_{1}\Sigma^{\prime}$ by putting the base point at the center of the circle and representing the generator $c_{i}$ by a path that runs outward along a radius of the circle, loops counterclockwise around $i$, and then returns along a radius to the center. More generally, for any set $I$ of consecutive punctures, let $c_{I}$ be the element of $\pi_{1}\Sigma^{\prime}$ represented by a path that runs outward along a radius, loops counterclockwise around $I$, and then returns along a radius to the center.
@@
When $I$ is the whole set of punctures, there are many paths that fit this description, but they all represent the identity in $\pi_{1}\Sigma^{\prime}$. If $I$ and $J$ are adjacent, disjoint sets of consecutive punctures, with $I$ counterclockwise of $J$, then $c_{I}c_{J}=c_{I\cup J}$.
The problem of finding the desired pants decomposition for $\rho\colon\pi_{1}\Sigma^{\prime}\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ can be phrased as a combinatorial game. Before we explain the rules, we need to paint $\operatorname{\mathbb{H}}$ so that every point is a different color. We then give each puncture $i$ the color of the fixed point of $\rho(c_{i})$. More generally, we give each set of punctures $I$ the color of the fixed point of $\rho(c_{I})$. We know that $\rho(c_{I})$ is elliptic because DT representations are totally elliptic, and $c_{I}$ is represented by a simple curve. Observation (2) gives this coloring a useful combinatorial property.
@@
By switching the labels of the sets if necessary, we can assume that $I$ is counterclockwise of $J$. We observed earlier that $c_{I}c_{J}=c_{I\cup J}$, implying that $\rho(c_{I}),\rho(c_{J})$, and $\rho(c_{I\cup J})^{-1}$ are three elliptic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ that can be multiplied to get the identity. Observation (2) then gives the desired conclusion. ∎
@@
For every DT representation, there exists a geometric presentation of $\pi_{1}\Sigma^{\prime}$ such that every triangle in the triangle chain associated to its standard pants decomposition has positive area.
@@
  • In the second move of the game, we removed either puncture $i-1$ or $i+2$. If we removed puncture $i+2$, then we simply define $c_{3}^{\prime}=c_{i+2}$. The second pants curve is then represented by the fundamental group element $c_{\{i,i+1,i+2\}}=c_{i}c_{i+1}c_{i+2}$. In this case we also define $b_{2}=(c_{\{i,i+1,i+2\}})^{-1}$. If instead we removed puncture $i-1$, then we alternatively define $c_{3}^{\prime}=b_{1}c_{i-1}b_{1}^{-1}$. In this case, the second pants curve is represented by $c_{\{i-1,i,i+1\}}=c_{i-1}c_{i}c_{i+1}$ and we define $b_{2}=(c_{\{i-1,i,i+1\}})^{-1}$. Observe that, in both cases, it holds that
@@
  • Now assume we have just defined $c_{j}^{\prime}$ with $33\leq j\leq n-$ 3 by modifying the fundamental group generator that we removed on the $(j-1)$th move. Let $I$ be the set of missing punctures at the beginning of the $j$th move and $z$ be the puncture removed on the $j$th move. Part of our induction hypotheses is that the last pants curve that we built is represented by the fundamental group element $b_{j-1}=c_{I}^{-1}=(c_{j}^{\prime})^{-1}\cdots(c_{2}^{\prime})^{-1}(c_{1}^{\prime})^{-1}$. If $z$ was to the right of $I$, then we let $c_{j+1}^{\prime}=c_{z}$. If $z$ was to the left of $I$, then we let $c_{j+1}^{\prime}=b_{j-1}c_{z}b_{j-1}^{-1}$. In both cases, the new pants curve is represented by
@@
  • Once we have defined $c_{n-2}^{\prime}$, then we finish the construction as follows. The last pant curve that was built is represented by the fundamental group element
@@
    where $I$ is now a set of all but two neighbouring punctures which we denote $x$ and $y$. Assume that $x$ is directly to the right of $I$. We define $c_{n-1}^{\prime}=c_{x}$ and $c_{n}^{\prime}=c_{y}$. Since $x$ is directly to the right of $I$ and $c_{1}\cdots c_{n}=1$, it holds that $(c_{I})^{-1}=c_{x}c_{y}=c_{n-1}^{\prime}c_{n}^{\prime}$. Comparing with the expression of $b_{n-3}$ above, we conclude that $c_{1}^{\prime}\cdots c_{n}^{\prime}=1$.
We have produced a collection $c_{1}^{\prime},\ldots,c_{n}^{\prime}$ of fundamental group element satisfying $c_{1}^{\prime}\cdots c_{n}^{\prime}=1$. By construction, each of the $c_{j}^{\prime}$ is conjugate to one of the generators $c_{1},\ldots,c_{n}$. Moreover, the pants curves of $\mathcal{B}$ are represented by $b_{j}=(c_{j+1}^{\prime})^{-1}\cdots(c_{2}^{\prime})^{-1}(c_{1}^{\prime})^{-1}$ making $\mathcal{B}$ the standard pants decomposition of $\Sigma^{\prime}$ for the geometric generators $c_{1}^{\prime},\ldots,c_{n}^{\prime}$ of $\pi_{1}\Sigma^{\prime}$. ∎
We can even go one step further as Proposition 3 to make sure that adjacent triangles in a chain don’t overlap. Two triangles are said to overlap if their interiors aren’t disjoint. Note that non-overlapping triangles are allowed to intersect along their boundary. In terms of the action-angle coordinates of [Mar24], two non-degenerate adjacent triangles with common vertex corresponding to the pants curve $b$ overlap if and only if $\pi<\gamma_{b}<2\pi$ when $\gamma_{b}$ is seen as a number in $[0,2\pi)$. When $\gamma_{b}\in\{0,\pi\}$, the triangles touch along their boundary.
@@
For every DT representation $\rho\colon\pi_{1}\Sigma^{\prime}\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, there exists a geometric presentation of $\pi_{1}\Sigma^{\prime}$ such that every triangle in the triangle chain associated to its standard pants decomposition has positive area and no pair of adjacent triangles overlap. [^14] In fact, we’ll prove something slightly stronger, as discussed in Remark 6.
@@
We start from the geometric presentation $c_{1},\ldots,c_{n}$ provided by Proposition 3. It guarantees that all the triangles in the chain associated to $\rho$ and built using the standard pants decomposition $b_{1},\ldots,b_{n-3}$ are non-degenerate. Assume that the two triangles that share the common vertex corresponding to the fixed point of $\rho(b_{i})$ overlap. In terms of action-angle coordinates, this means that $\gamma_{i}\in(\pi,2\pi)$.
Consider the fundamental group automorphism $\tau_{i}\colon\pi_{1}\Sigma^{\prime}\to\pi_{1}\Sigma^{\prime}$ defined by $\tau_{i}(c_{j})=c_{j}$ if $j\leq i+1$ and $\tau_{i}(c_{j})=b_{i}c_{j}b_{i}^{-1}$ if $j\geq i+2$. It’s not hard to see that $\tau_{i}(c_{1})\cdots\tau_{i}(c_{n})=1$ using that $b_{i}=c_{i+1}^{-1}c_{i}^{-1}\cdots c_{1}^{-1}$. (Under the classical Dehn-Nielsen-Baer correspondence, explained for instance in [FM12, §8], the automoprhism $\tau_{i}$ comes from the Dehn twist along the simple closed curve $b_{i}$.) Changing the generating family of $\pi_{1}\Sigma^{\prime}$ from $\{c_{1},\ldots,c_{n}\}$ to $\tau_{i}(\{c_{1},\ldots,c_{n}\})$ only affects one action-angle coordinate of $\rho$: the angle coordinate $\gamma_{i}$. The change in $\gamma_{i}$ is given by the formula

### v3

@@ 4.4.2. Injectivity of the realization map
The value of $\cos(\beta_{b})$ is determined by (4), and the value of $\beta_{b}$ is then determined by the sign of $\kappa_{b}$:
@@
Next, we prove that $[\mathcal{F}]$ also determines the slit angles. We’re assuming that the underlying pants assembly $\Upsilon$ is chained, so there are never more than two slits on a given samosa. That means we can find expressions for all the slit angles using Lemma 3:
@@ 4.4.8. Continuity of the inverse map
We now prove that the inverse map of the realization map, defined on $R(\mathcal{S}_{\alpha,\Upsilon}^{\varepsilon})\subset\operatorname{Hyp}_{\alpha}(S_{\Upsilon})$, is also continuous. To do so, we show that the samosa assembly parameters $(\beta_{b},\ell_{b},\phi_{b},\phi^{\prime}_{b})_{b\in{\mathcal{B}}}$ of $R^{-1}([\mathcal{F}])$ depend continuously on $[\mathcal{F}]\in R(\mathcal{S}_{\alpha,\Upsilon}^{\varepsilon})$. We can relate $\ell_{b}$ to $c_{b}$ and $\kappa_{b}$ by (10). Proposition 1 says that $c_{b}$ depends continuously on $[\mathcal{F}]$. We can use hyperbolic trigonometry to express the kink angle $\kappa_{b}$ in terms of triangulation edge lengths, so Proposition 1 implies that it too depends continuously on $[\mathcal{F}]$. This shows that $\ell_{b}$ depends continuously on $[\mathcal{F}]$ for every $b\in{\mathcal{B}}$. The value of $\beta_{b}$ is expressed continuously in terms of $c_{b}$ and $\kappa_{b}$ by (11), showing that $\beta_{b}$ varies continuously in $[\mathcal{F}]$.
Finally, to prove that the slit angles $\phi_{b}$ and $\phi_{b}^{\prime}$ depend continuously on $[\mathcal{F}]$, we use (12) and (13). These equations express the slit angles as continuous functions of $\alpha$, $\beta$, $\ell$, $\lambda$, and sometimes $\xi$ and $\eta$. We already established the continuous dependence of $\ell_{b}$ and $\beta_{b}$ on $[\mathcal{F}]$. From Proposition 1, we deduce that both $\lambda$ and $\xi$ depend continuously on $[\mathcal{F}]$. Since $\eta$ is an angle of the triangulation, the hyperbolic law of cosines—formula (C.1) from Appendix A—shows that it too depends continuously on $[\mathcal{F}]$. This finishes the proof that $R^{-1}$ is continuous.
@@ 4.4.9. Smoothness of the realization map and its inverse
It remains to be shown that if Conjecture 1 holds, the realization map is a diffeomorphism onto its image. The argument is similar to the proof that the realization map is continuous.
Using Conjecture 1, we can prove that $R$ is smooth by showing that the edge lengths of various triangulations depend smoothly on the samosa parameters. We use (14), (15), (16), (17) for that. The non-smoothness of (16) at $\phi_{b}=0$ does no harm because we’re working on the domain $\mathcal{S}_{\alpha,\Upsilon}^{\varepsilon}$, which prescribes a hemisphere for each slit, and hence a sign for each slit angle.
To prove that the inverse map is smooth, it’s enough to express the samosa parameters as smooth functions of the same triangulation edge lengths. Here, we use (10), (11), (12), (13). For instance, we see from (11) that $\beta_{b}$ depends smoothly on the parameters $\kappa_{b}$ and $c_{b}$, except maybe when $\kappa_{b}=0$ or when
@@
Using trigonometric identities, we can see that $\sin(\kappa_{b}/2)^{2}\cosh(c_{b})-\cos(\kappa_{b}/2)^{2}\geq-1$, with equality if and only if $\kappa_{b}=0$. To investigate the situation at $\kappa_{b}=0$, we can write
@@
using the hyperbolic law of sines—formula (S) from Appendix A. This alternate expression for $\beta_{b}$ shows that it depends smoothly on $\kappa_{b}$, $c_{b}$ and $\ell_{b}$ around $\kappa_{b}=0$ too, because $c_{b}$ and $\ell_{b}$ stay bounded away from zero in this region. The treatment of the other variables is similar.
@@ 5.2. Unfolding samosas
A samosa with two slits in the same hemisphere is a hamantash as long as its slit angles satisfy $\phi_{a}\neq 0$ and $\phi_{b}^{\prime}\neq\pm\beta_{b}^{\prime}/2$, and its slit lengths $\ell_{a}$ and $\ell_{b}$ don’t exceed the respective bounds $\ell_{a,\textup{max}}$ and $\ell_{b,\textup{max}}$ given by the formulas
$\coth(\ell_{a,\textup{max}})$ $=\frac{1}{\sinh(d)}\big{(}\cos(\phi_{a})\cosh(d)+|\sin(\phi_{a})|\cot(\beta_{b}^{\prime}/2-|\phi_{b}^{\prime}|)\big{)}$
$\coth(\ell_{b,\textup{max}})$ $=\frac{1}{\sinh(d)}\big{(}\cos(\beta_{b}^{\prime}/2-|\phi_{b}^{\prime}|)\cosh(d)+\sin(\beta_{b}^{\prime}/2-|\phi_{b}^{\prime}|)\cot(|\phi_{a}|)\big{)},$
@@ 5.3. Unfolding samosa assemblies
The subspace of hamantash assemblies is denoted $\mathcal{H}_{\alpha,\Upsilon}\subset\mathcal{S}_{\alpha,\Upsilon}$. A hamantash assembly can be unfolded to a collection of hyperbolic polygons embedded in the hyperbolic plane. Using side identifications, all the pieces can be glued back together to produce the realization of the hamantash assembly. The process of gluing the hamantashen along their slits commutes with the processes of unfolding and re-folding the hamantashen. In particular, we can glue the hamantashen along their slits while they’re laid out, unfolded, on the hyperbolic plane. This produces a net for the realization—although the net might overlap itself.
When none of the slits lie on equator segments opposite from un-slit corners, as discussed in Section 4.3.4, the realized hamantash assembly comes with a canonical isotopy class of homeorphisms to the generic realization $S_{\Upsilon}$. Thus, the process of unfolding, gluing along the slits, and then re-folding is the restriction $R\colon\mathcal{H}_{\alpha,\Upsilon}^{\text{gen}}\to\operatorname{Hyp}_{\alpha}(S_{\Upsilon})$ of the realization map, where $\mathcal{H}_{\alpha,\Upsilon}^{\text{gen}}=\mathcal{S}_{\alpha,\Upsilon}^{\text{gen}}\cap\mathcal{H}_{\alpha,\Upsilon}$.
@@ 6.1. The holonomies of hamantashen
Given a chained pants assembly $\Upsilon$, let $\mathcal{H}_{\alpha,\Upsilon}^{\textrm{north}}\subset\mathcal{H}_{\alpha,\Upsilon}$ be the subset consisting of hamantash assemblies whose slits all lie in northern hemispheres. The realization map extends continuously from $\mathcal{H}_{\alpha,\Upsilon}^{\text{gen}}$ to $\mathcal{H}_{\alpha,\Upsilon}^{\text{north}}$. To see why, first recall from Section 4.3.4 why the realization map fails to extend from $\mathcal{S}_{\alpha,\Upsilon}^{\text{gen}}$ to $\mathcal{S}_{\alpha,\Upsilon}$. The problem is that a non-generic samosa assembly comes with several isometry classes of identifications with $S_{\Upsilon}$, and the choice between them can’t be made continuously throughout $\mathcal{S}_{\alpha,\Upsilon}$. In $\mathcal{H}_{\alpha,\Upsilon}^{\text{north}}$, we can make a continuous choice of identification by pushing all the slits infinitesimally into the interior of the northern hemisphere.
@@
can be understood as follows. Unfolding a hamantash assembly produces a collection of hyperbolic polygons embedded in the hyperbolic plane. These polygons are glued together as described after Definition 9. An unfolded hamantash in $\mathcal{H}_{\alpha,\Upsilon}^{\textrm{north}}$ consists of a triangle, which is the southern hemisphere, and one to three flaps, which are the pieces of the northern hemisphere. For every pants curve in $\mathcal{B}$, there are two triangles meeting at a shared vertex, with the adjacent flaps glued together along the cut edges of a slit. Erasing the flaps gives a triangle chain, laid out along the spine of the hamantash assembly, with the geometric features described in Section 2.2. Since all the triangles in the chain are southern hemispheres, they all lie on the same side of the spine introduced in Section 4.3.4.
@@ 6.2. Finding a suitable pants decomposition
    For any pair of pants decomposition and any DT representation, there is always at least one non-degenerate triangle in the associated chain. This follows from the fact that the total area of the triangles in the chain is a positive constant (essentially equal to the Toledo number of the representation).
@@
Assume that $n\geq 4$, since the claim is obvious for $n=3$. For concreteness, identify $\Sigma$ with the one-point compactification of the plane, with the marked points laid out around a circle. Label the marked points $1,\ldots,n$, going counterclockwise around the circle. The marked points then form a cyclically ordered set. Introduce a standard geometric presentation of $\pi_{1}\Sigma^{\prime}$ by putting the base point at the center of the circle and representing the generator $c_{i}$ by a path that runs outward along a radius of the circle, loops counterclockwise around $i$, and then returns along a radius to the center. More generally, for any set $I$ of consecutive punctures, let $c_{I}$ be the element of $\pi_{1}\Sigma^{\prime}$ represented by a path that runs outward along a radius, loops counterclockwise around $I$, and then returns along a radius to the center.
@@
When $I$ is the whole set of punctures, there are many paths that fit this description, but they all represent the identity in $\pi_{1}\Sigma^{\prime}$. If $I$ and $J$ are adjacent, disjoint sets of consecutive punctures, with $I$ clockwise of $J$, then $c_{I}c_{J}=c_{I\cup J}$. [^15] In fundamental group product expressions, we concatenate the loops from left to right, so $c_{I}c_{J}$ means following $c_{I}$ and then $c_{J}$.
The problem of finding the desired pants decomposition for $\rho\colon\pi_{1}\Sigma^{\prime}\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ can be phrased as a combinatorial game. Before we explain the rules, we need to paint $\operatorname{\mathbb{H}}$ so that every point is a different color. We then give each puncture $i$ the color of the fixed point of $\rho(c_{i})$. More generally, we give each set of consecutive punctures $I$ the color of the fixed point of $\rho(c_{I})$. We know that $\rho(c_{I})$ is elliptic because DT representations are totally elliptic, and $c_{I}$ is represented by a simple curve. Observation (2) gives this coloring a useful combinatorial property.
@@
By switching the labels of the sets if necessary, we can assume that $I$ is clockwise of $J$. We observed earlier that $c_{I}c_{J}=c_{I\cup J}$, implying that $\rho(c_{I}),\rho(c_{J})$, and $\rho(c_{I\cup J})^{-1}$ are three elliptic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ that can be multiplied to get the identity. Observation (2) then gives the desired conclusion. ∎
@@
For every DT representation, there exists a geometric presentation of $\pi_{1}\Sigma^{\prime}$ such that every triangle in the triangle chain associated with its standard pants decomposition has positive area.
@@
  • In the second move of the game, we removed either puncture $i-1$ or $i+2$. If we removed puncture $i+2$, then we simply define $c_{3}^{\prime}=c_{i+2}$. The second pants curve is then represented by the fundamental group element $c_{\{i,i+1,i+2\}}=c_{i}c_{i+1}c_{i+2}$. In this case we also define $b_{2}=(c_{\{i,i+1,i+2\}})^{-1}$. If instead we removed puncture $i-1$, then we alternatively define $c_{3}^{\prime}=b_{1}c_{i-1}b_{1}^{-1}$. In this case, the second pants curve is represented by $c_{\{i-1,i,i+1\}}=c_{i-1}c_{i}c_{i+1}$ and we define $b_{2}=(c_{\{i-1,i,i+1\}})^{-1}$. In both cases,
@@
  • Now, suppose we’ve just defined $c_{j}^{\prime}$ with $33\leq j\leq n-$ 3 by modifying the fundamental group generator that we removed on the $(j-1)$th move. Let $I$ be the set of missing punctures at the beginning of the $j$th move and $z$ be the puncture removed on the $j$th move. Part of our induction hypotheses is that the last pants curve that we built is represented by the fundamental group element $b_{j-1}=c_{I}^{-1}=(c_{j}^{\prime})^{-1}\cdots(c_{2}^{\prime})^{-1}(c_{1}^{\prime})^{-1}$. If $z$ was directly counterclockwise of $I$, then we let $c_{j+1}^{\prime}=c_{z}$. If $z$ was clockwise of $I$, then we let $c_{j+1}^{\prime}=b_{j-1}c_{z}b_{j-1}^{-1}$. In both cases, the new pants curve is represented by
@@
  • Once we’ve defined $c_{n-2}^{\prime}$, then we finish the construction as follows. The last pants curve we built is represented by the fundamental group element
@@
    where $I$ is now a set of all but two neighbouring punctures which we denote $x$ and $y$. Assume that $x$ is directly counterclockwise of $I$. We define $c_{n-1}^{\prime}=c_{x}$ and $c_{n}^{\prime}=c_{y}$. Since $x$ is directly counterclockwise of $I$ and $c_{1}\cdots c_{n}=1$, it holds that $(c_{I})^{-1}=c_{x}c_{y}=c_{n-1}^{\prime}c_{n}^{\prime}$. Comparing with the expression of $b_{n-3}$ above, we conclude that $c_{1}^{\prime}\cdots c_{n}^{\prime}=1$.
We have produced a collection $c_{1}^{\prime},\ldots,c_{n}^{\prime}$ of fundamental group element satisfying $c_{1}^{\prime}\cdots c_{n}^{\prime}=1$. By construction, each of the $c_{j}^{\prime}$ is conjugate to one of the generators $c_{1},\ldots,c_{n}$. Moreover, the pants curves of $\mathcal{B}$ are represented by $b_{j}=(c_{j+1}^{\prime})^{-1}\cdots(c_{2}^{\prime})^{-1}(c_{1}^{\prime})^{-1}$, making $\mathcal{B}$ the standard pants decomposition of $\Sigma^{\prime}$ for the geometric generators $c_{1}^{\prime},\ldots,c_{n}^{\prime}$ of $\pi_{1}\Sigma^{\prime}$. ∎
We can even go one step further than Proposition 3 to make sure that adjacent triangles in a chain don’t overlap. Two triangles are said to overlap if their interiors aren’t disjoint. Note that non-overlapping triangles are allowed to intersect along their boundaries. In terms of the action-angle coordinates of [Mar24], two non-degenerate adjacent triangles with common vertex corresponding to the pants curve $b$ overlap if and only if $\pi<\gamma_{b}<2\pi$ when $\gamma_{b}$ is seen as a number in $[0,2\pi)$. When $\gamma_{b}\in\{0,\pi\}$, the triangles touch along their boundaries.
@@
For every DT representation $\rho\colon\pi_{1}\Sigma^{\prime}\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, there exists a geometric presentation of $\pi_{1}\Sigma^{\prime}$ such that every triangle in the triangle chain associated with its standard pants decomposition has positive area and no pair of adjacent triangles overlap. [^16] In fact, we’ll prove something slightly stronger, as discussed in Remark 6.
@@
We start from the geometric presentation $c_{1},\ldots,c_{n}$ provided by Proposition 3. It guarantees that all the triangles in the chain associated to $\rho$ and built using the standard pants decomposition $b_{1},\ldots,b_{n-3}$ are non-degenerate. Suppose that the two triangles that share the vertex corresponding to the fixed point of $\rho(b_{i})$ overlap. In terms of action-angle coordinates, this means that $\gamma_{i}\in(\pi,2\pi)$.
Consider the fundamental group automorphism $\tau_{i}\colon\pi_{1}\Sigma^{\prime}\to\pi_{1}\Sigma^{\prime}$ defined by
$\tau_{i}(c_{j})=\begin{cases}c_{j}&j\leq i+1\\ b_{i}c_{j}b_{i}^{-1}&j\geq i+2.\end{cases}$
We can see that $\tau_{i}(c_{1})\cdots\tau_{i}(c_{n})=1$ by unpacking the definition of $\tau_{i}$ and recalling that $b_{i}=c_{i+1}^{-1}c_{i}^{-1}\cdots c_{1}^{-1}$. (Under the classical Dehn–Nielsen–Baer correspondence, explained for instance in [FM12, §8], the automoprhism $\tau_{i}$ comes from the Dehn twist along the simple closed curve $b_{i}$.) Changing the generating family of $\pi_{1}\Sigma^{\prime}$ from $\{c_{1},\ldots,c_{n}\}$ to $\tau_{i}(\{c_{1},\ldots,c_{n}\})$ only affects one action-angle coordinate of $\rho$: the angle coordinate $\gamma_{i}$. The change in $\gamma_{i}$ is given by the formula